### v2

# Neuronal Attention Circuit (NAC) for Representation Learning

Abstract
Attention improves representation learning over RNNs, but its discrete nature limits continuous-time (CT) modeling. We introduce Neuronal Attention Circuit (NAC), a novel, biologically plausible CT-Attention mechanism that reformulates attention logits computation as the solution to a linear first-order ODE with nonlinear interlinked gates derived from repurposing C. elegans Neuronal Circuit Policies (NCPs) wiring mechanism. NAC replaces dense projections with sparse sensory gates for key-query projections and a sparse backbone network with two heads for computing content-target and learnable time-constant gates, enabling efficient adaptive dynamics. NAC supports three attention logit computation modes: (i) explicit Euler integration, (ii) exact closed-form solution, and (iii) steady-state approximation. To improve memory intensity, we implemented a sparse Top-K pairwise concatenation scheme that selectively curates key-query interactions. We provide rigorous theoretical guarantees, including state stability, bounded approximation errors, and universal approximation. Empirically, we implemented NAC in diverse domains, including irregular time-series classification, lane-keeping for autonomous vehicles, and industrial prognostics. We observed that NAC matches or outperforms competing baselines in accuracy and occupies an intermediate position in runtime and memory efficiency compared with several CT baselines.
Continuous-time Attention, Liquid Neural Networks (LNNs)
1 Introduction
Learning representations of sequential data in temporal or spatio-temporal domains is essential for capturing patterns and enabling accurate forecasting. Discrete-time Recurrent neural networks (DT-RNNs) such as RNN (rumelhart1985learning; jordan1997serial), Long-short term memory (LSTM) (hochreiter1997long), and Gated Recurrent Unit (GRU) (cho2014learning) model sequential dependencies by iteratively updating hidden states to represent or predict future elements in a sequence. While effective for regularly sampled sequences, DT-RNNs face challenges with irregularly sampled data because they assume uniform time intervals. In addition, vanishing gradients can make it difficult to capture long-term dependencies (hochreiter1998vanishing).
Continuous-time RNNs (CT-RNNs) (rubanova2019latent) model hidden states as ordinary differential equations (ODEs), allowing them to process inputs that arrive at arbitrary or irregular time intervals. Mixed-memory RNNs (mmRNNs) (lechner2022mixed) build on this idea by separating memory compartments from time-continuous states, helping maintain stable error propagation while capturing continuous-time dynamics. Liquid neural networks (LNNs) (hasani2021liquid; hasani2022closed) take a biologically inspired approach by assigning variable time-constants to hidden states, improving adaptability and robustness, though vanishing gradients can still pose challenges during training.
The attention mechanisms (vaswani2017attention) mitigate this limitation by treating all time steps equally and allowing models to focus on the most relevant observations. It computes the similarity between queries ($q$) and keys ($k$), scaling by the key dimension to keep gradients stable. Multi-Head Attention (MHA) (vaswani2017attention) extends this by allowing the model to attend to different representation subspaces in parallel. Variants like Sparse Attention (tay2020sparse; roy2021efficient), BigBird (zaheer2020big), and Longformer (beltagy2020longformer) modify the attention pattern to reduce computational cost, particularly for long sequences, by attending only to selected positions rather than all pairs. Even with these improvements, attention-based methods still rely on discrete scaled dot-product operations, limiting their ability to model continuous trajectories often captured by CT counterparts.
Recent work has explored bridging this gap through Neural-ODE (chen2018neural) formulation. mTAN (shukla2021multi) learns CT embeddings and uses time-based attention to interpolate irregular observations into a fixed-length representation for downstream encoder-decoder modeling. ODEFormer (d2023odeformer) trains a sequence-to-sequence transformer on synthetic trajectories to directly infer symbolic ODE systems from noisy, irregular data, though it struggles with chaotic systems and generalization beyond observed conditions. Continuous-time Attention (CTA) (chien2021continuous) embeds a continuous-time attention mechanism within a Neural ODE, allowing attention weights and hidden states to evolve jointly over time. Still, it remains computationally intensive and sensitive to the accuracy of the ODE solver. ContiFormer (chen2023contiformer) builds a CT-transformer by pairing ODE-defined latent trajectories with a time-aware attention mechanism to model dynamic relationships in data.
Despite these innovations, a persistent and underexplored gap remains in developing a biologically plausible attention mechanism that seamlessly integrates CT dynamics with the abstraction of the brain’s connectome to handle irregular sequences without prohibitive computational costs. Building on this, we propose a novel attention mechanism called the Neuronal Attention Circuit (NAC), in which attention logits are computed as the solution to a first-order ODE modulated by nonlinear, interlinked gates derived from repurposing Neuronal Circuit Policies (NCPs) from the nervous system of C. elegans nematode (refer to Appendix A.2 for more information). Unlike standard attention, which projects key-query pairs through a dense layer, NAC employs a sensory gate to transform input features and a backbone to model nonlinear interactions, with multiple heads producing outputs structured for attention logits computation. Based on the solutions to ODE, we define three computation modes: (i) Exact, using the closed-form ODE solution; (ii) Euler, approximating the solution via explicit Euler integration; and (iii) Steady, using only the steady-state solution, analogous to standard attention scores. To reduce computational complexity, we implemented a sparse Top-K pairwise concatenation algorithm that selectively curates key-query inputs. We evaluate NAC across multiple domains, including irregularly sampled time series, autonomous vehicle lane-keeping, and Industry 4.0, comparing it to state-of-the-art baselines. NAC consistently matches or outperforms these models, while runtime and peak memory benchmarks place it between CT-RNNs in terms of speed and CT-Attentions in terms of memory requirements.
2 Neuronal Attention Circuit (NAC)
We propose a simple alternative formulation of the attention logits $a$ (refer to Appendix A.1 for more information), interpreting them as the solution to a first-order linear ODE modulated by nonlinear, interlinked gates:
$\frac{da_{t}}{dt}=-\,\underbrace{f_{\omega_{\tau}}([\mathbf{q};\mathbf{k}],\theta_{\omega_{\tau}})}_{\omega_{\tau}(\mathbf{u})}a_{t}+\underbrace{f_{\phi}([\mathbf{q};\mathbf{k}],\theta_{\phi})}_{\phi(\mathbf{u})},$ (1)
where $\mathbf{u}=[\mathbf{q};\mathbf{k}]$ denotes the sparse Top-K concatenated query–key input. $\omega_{\tau}$ represents a learnable time-constant gate head, $\phi$ denotes a nonlinear content-target head. Both gates are parameterized by a backbone network derived from repurposing NCPs. We refer to this formulation as the Neuronal Attention Circuit (NAC). It enables the logits $a_{t}$ to evolve dynamically with input-dependent, variable time constants, mirroring the adaptive temporal dynamics found in C. elegans nervous systems while improving computational efficiency and expressiveness. Moreover, it introduces continuous depth into the attention mechanism, bridging discrete-layer computation with dynamic temporal evolution.
Motivation behind this formulation: The proposed formulation is loosely motivated by the input-dependent time-constant mechanism of Liquid Neural Networks (LNNs), a class of CT-RNNs inspired by biological nervous systems and synaptic transmission. In this framework, the dynamics of non-spiking neurons are described by a linear ODE with nonlinear, interlinked gates: $\frac{d\mathbf{x_{t}}}{dt}=\frac{\mathbf{x_{t}}}{\tau}+\mathbf{S_{t}},$ where $\mathbf{x_{t}}$ denotes the hidden state and $\mathbf{S_{t}}\in\mathbb{R}^{M}$ represents a nonlinear contribution defined as $f(\mathbf{x_{t}},\mathbf{u},t,\theta)(A-\mathbf{x_{t}})$. Here, $A$ and $\theta$ are learnable parameters. Plugging $\mathbf{S_{t}}$ yields $\frac{d\mathbf{x_{t}}}{dt}=-\left[-\frac{1}{\tau}-f(\mathbf{x_{t}},\mathbf{u},t,\theta)\right]\mathbf{x_{t}}+f(\mathbf{x_{t}},\mathbf{u},t,\theta)A$. LNNs are known for their strong expressivity, stability, and performance in irregularly sampled time-series modeling (hasani2021liquid; hasani2022closed).
NAC’s forward-pass update using ODE solver: The state of NAC at time $t$ can be computed using a numerical ODE solver that simulates the dynamics from an initial state $a_{0}$ to $a_{t}$. The solver discretizes the continuous interval $[0,T]$ into steps $t_{0},t_{1},t_{2},\dots,t_{n}]$, with each step updating the state from $t_{i}$ to $t_{i+1}$. For our purposes, we use the explicit Euler solver, which is simple, efficient, and easy to implement. Although methods such as Runge-Kutta may offer higher accuracy, their computational overhead makes them less suitable for large-scale neural simulations that require numerous updates, especially since the logits are normalized, and exact precision is not necessary. Let the step size be $\Delta t$, with discrete times $t_{n}=n\Delta t$ and logit states $a_{n}=a(t_{n})$. Using the explicit Euler method, the update is
$a_{n+1}=a_{n}+\Delta t\,(-\omega_{\tau}a_{n}+\phi).$ (2)
Closed-form (Exact) Computation of NAC: We now devise the analytical solution for Eqn. 1. Let both $\omega_{\tau}$ and $\phi$ be fixed in pseudo-time interval (frozen-coefficient approximation (john1952integration)) with initial condition $a_{0}$, then closed-form solution is:
$a_{t}=\underbrace{a^{*}}_{\text{steady-state}}+\underbrace{(a_{0}-a^{*})\,e^{-\omega_{\tau}t}}_{\text{transient}}$ (3)
Here, $a^{*}=\phi/\omega_{\tau}$ is the steady-state solution. The full derivation is provided in Appendix B.1.
2.1 Stability Analysis of NAC
We now investigate the stability bounds of NAC under both the ODE-based and the Closed-Form formulations.
2.1.1 State Stability
We analyze state stability in both single-connection and multi-connection settings. This analysis establishes the boundedness of the attention logit state trajectory, ensuring that, under positive decay rates, the dynamics remain well-behaved without divergence or overshoot.
Theorem 1 (State Stability).
Let $a_{t}^{(i)}$ denote the state of the $i$-th attention logit governed by $da_{t}^{(i)}/dt=-\omega_{\tau}a_{t}^{(i)}+\phi$. Assume that $\phi$ and $\omega_{\tau}$ decompose across $M$ incoming connections as $\phi=\sum_{j=1}^{M}f_{\phi}([\mathbf{q}_{i};\mathbf{k}_{j}])$, and $\omega_{\tau}=\sum_{j=1}^{M}f_{\omega_{\tau}}([\mathbf{q}_{i};\mathbf{k}_{j}]),$ with $f_{\omega_{\tau}}>0$. Define the per-connection equilibrium $A_{i,j}=f_{\phi}([\mathbf{q}_{i};\mathbf{k}_{j}])/f_{\omega_{\tau}}([\mathbf{q}_{i};\mathbf{k}_{j}])$, and let $A_{i}^{\min}=\min_{j}A_{i,j}$ and $A_{i}^{\max}=\max_{j}A_{i,j}$. Then for any finite horizon $t\in[0,T]$, the state trajectory satisfies
$\min(0,A_{i}^{\min})\;\leq\;a_{t}^{(i)}\;\leq\;\max(0,A_{i}^{\max}),$ (4)
provided the initial condition $a_{i}(0)$ lies within this range. In the special case of a single connection ($M=1$), the bounds collapse to
$\min(0,A_{i})\;\leq\;a_{t}^{(i)}\;\leq\;\max(0,A_{i}),$ (5)
where $A_{i}=f_{\phi}/\omega_{\tau}$ is exactly the steady-state solution from Eqn. 3. The proof is provided in the Appendix B.2.
2.1.2 Closed-form Error & Exponential Bounds
We now examine the asymptotic stability, error characterization, and exponential boundedness of the closed-form formulation. We begin by quantifying the deviation of the trajectory from its steady-state solution. Define the instantaneous error
$\varepsilon_{t}=a_{t}-a^{*},$ (6)
which measures the distance of the system state to equilibrium at time $t$. From Eqn. 3, the error admits the exact representation
$\varepsilon_{t}=(a_{0}-a^{*})e^{-\omega_{\tau}t}$ (7)
In particular, the pointwise absolute error is given by
$|\varepsilon_{t}|=|a_{0}-a^{*}|\,e^{-\omega_{\tau}t}$ (8)
This reveals that convergence is not merely asymptotic but follows an exact exponential law, controlled by the rate parameter $\omega_{\tau}$. This yields the following finite-time guarantee.
Corollary 1 (Exponential decay bound).
If $\omega_{\tau}>0$, then for all $t\geq 0$,
$|a_{t}-a^{*}|\leq|a_{0}-a^{*}|\,e^{-\omega_{\tau}t}.$ (9)
Remark: If $\omega_{\tau}>0$ then $\lim_{t\to\infty}e^{-\omega_{\tau}t}=0$, therefore $a_{t}\to a^{*}$. The convergence is exponential with rate $\omega_{\tau}$. If $\omega_{\tau}<0$ then $e^{-\omega_{\tau}t}=e^{|\omega_{\tau}|t}$ diverges so $a_{t}$ grows exponentially away from $a^{*}$ in magnitude (unless initial offset $a_{0}-a^{*}=0$, a measure-zero case). If $\omega_{\tau}=0$ the ODE is $\dot{a}=\phi$ and the solution is linear in $t$ (unless $\phi=0$). For bounded dynamics that converge to an interpretable steady-state, it is required that $\omega_{\tau}>0$.
Corollary 2 (Uniform initialization).
If the initialization is only known to belong to a bounded set, i.e., $|a_{0}-a^{*}|\leq M$ for some $M>0$, then the error admits the uniform bound
$|a_{t}-a^{*}|\leq Me^{-\omega_{\tau}t}.$ (10)
Remark: This bound highlights that exponential convergence holds uniformly across all admissible initial conditions, with the constant $M$ capturing the worst-case deviation.
Corollary 3 (Sample complexity to $\delta$-accuracy).
A natural operational question is the time required to achieve a target tolerance $\delta>0$. Solving
$|a_{0}-a^{*}|e^{-\omega_{\tau}t}\leq\delta,$ (11)
We obtain the threshold
$t\geq\frac{1}{\omega_{\tau}}\ln\frac{|a_{0}-a^{*}|}{\delta}.$ (12)
Remark: The convergence rate is inversely proportional to $\omega_{\tau}$, and the required time scales only logarithmically in the accuracy level $1/\delta$. Intuitively, larger $\omega_{\tau}$ accelerates contraction towards equilibrium, yielding faster attainment of any prescribed tolerance.
Figure 1: Illustration of (a) NCPs with pre-determined wiring; (b) Sensory gate, where sensory neurons are active, and the remaining neurons are disabled for the $q$, $k$, and $v$ projections; (c) Backbone, showing inter-motor projections with sensory neurons disabled in extended heads for computing $\phi$ and $\omega_{\tau}$.
Algorithm 1 Repurposed NCPCell
0:  Wiring $\mathcal{W}$ with $(A_{\text{in}},A_{\text{rec}})$, groups $(N_{s},N_{i},N_{c},N_{m})$, activation $\alpha$, input group $G_{\text{input}}$, output group $G_{\text{output}}$, disabled groups $\mathcal{D}$
0:  Output $y_{t}\in\mathbb{R}^{B\times d_{\text{out}}}$, state $\mathbf{x}_{t}\in\mathbb{R}^{B\times d_{h}}$
  Binary mask: $M_{\text{rec}}\leftarrow|A_{\text{rec}}|$,    $M_{\text{in}}\leftarrow|A_{\text{in}}|$
  Initialize parameters : $W_{\text{in}},W_{\text{rec}},b,$ $w_{\text{in}},b_{\text{in}},w_{\text{out}},b_{\text{out}}$
  Input neurons: $I_{\text{in}}\leftarrow G_{\text{input}}$
[…]
  Define activation mask: $\text{mask}_{\text{act},i}=0$ if $i\in\mathcal{D}$ else 11
  Input Projections: $\tilde{u}_{t}\leftarrow u_{t}\odot w_{\text{in}}+b_{\text{in}}$
  Recurrent computation: $r_{t}\leftarrow\mathbf{x}_{t-1}(W_{\text{rec}}\odot M_{\text{rec}})$
  Sparse computation: $s_{t}\leftarrow\tilde{u}_{t}(W_{\text{in}}\odot M_{\text{in}})$
  Neuron update: $\mathbf{x}_{t}\leftarrow\alpha(r_{t}+s_{t}+b)\odot\text{mask}_{\text{act}}$
  Output mapping: $y_{t}\leftarrow(\mathbf{x}_{t}[I_{\text{out}}]\odot w_{\text{out}})+b_{\text{out}}$
  return $(y_{t},\mathbf{x}_{t})$
Algorithm 2 Sparse Top-K Pairwise Concatenation
0:  Keys $K\in\mathbb{R}^{B\times H\times T_{k}\times D}$, Top-K value $K$
0:  concatenated tensor $U\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times 2D}$
  Scores: $S\leftarrow Q\cdot K^{\top}$
  Effective Top-K: $K_{\text{eff}}\leftarrow\min(K,T_{k})$
  Indices: $I_{\text{topk}}\leftarrow\text{top\_k}(S,K_{\text{eff}})$
  Gather: $K_{\text{selected}}\leftarrow\text{gather}(K,I_{\text{topk}})\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times D}$
  Tiled: $Q_{\text{tiled}}\leftarrow\text{tile}(Q,K_{\text{eff}})\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times D}$
  Concatenate: $U_{\text{topk}}\leftarrow[\,Q_{\text{tiled}};\;K_{\text{selected}}\,]\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times 2D}$
  return $U_{\text{topk}}$
2.2 Designing the Neural Network
We now outline the design of a neural network layer guided by the preceding analysis. The process involves five steps: (i) repurposing NCPs; (ii) input curation; (iii) construction of the time vector ($t$); (iv) computing attention logits and weights; and (v) generating the attention output. Figure 2 provides a graphical overview of NAC.
Repurposing NCPs: We repurpose the NCPs framework by converting its fixed, biologically derived wiring (see Figure 1(a)) into a flexible recurrent architecture that allows configurable input–output mappings. Instead of enforcing a static connectome, our approach exposes adjacency matrices as modifiable structures defining sparse input and recurrent connections. This enables selective information routing across neuron groups while retaining the original circuit topology. Decoupling wiring specifications from model instantiation allows dynamic connectivity adjustments to accommodate different input modalities without full retraining. Algorithm 1 summarizes the steps for repurposing the NCPs wiring mechanism. Key features include group-wise masking for neuron isolation, adaptive remapping of inputs and outputs for task-specific adaptation, and tunable sparsity $s$ to balance expressiveness and efficiency.
In our implementation, the sensory neuron gate ($\mathcal{NN}_{\text{sensory}}$) projects the $q$, $k$, and $v$ representations (see Figure 1(b)). This enables sensory neurons to maintain structured, context-aware representations rather than collapsing inputs into fully connected layers. As a result, the network preserves locality and modularity, which improves information routing.
$\mathcal{NN}_{\text{sensory}}$ $=\text{NCPCell}(G_{\text{input}}=[N_{s}],\,G_{\text{output}}=[N_{s}],$ (13)
$\qquad\quad\mathcal{D}=[N_{i},N_{c},N_{m}],\,s)$
The inter-to-motor pathways form a backbone network ($\mathcal{NN}_{\text{backbone}}$) with branches that compute $\phi$ and $\omega_{\tau}$ (see Figure 1(c)). Instead of learning $\phi$ and $\omega_{\tau}$ independently, this backbone allows the model to learn shared representations, enabling multiple benefits: (i) separate head layers enable the system to capture temporal and structural dependencies independently; (ii) accelerates convergence during training.
$\mathcal{NN}_{\text{backbone}}$ $=\text{NCPCell}(G_{\text{input}}=[N_{i}],\,G_{\text{output}}=[N_{m}],$ (14)
$\qquad\quad\mathcal{D}=[N_{s}],\,s)$
The output heads are defined as:
$\phi$ $=\sigma(\mathcal{NN}_{\text{backbone}}(\mathbf{u}))$ (15)
$\omega_{\tau}$ $=\mathrm{softplus}(\mathcal{NN}_{\text{backbone}}(\mathbf{u}))+\varepsilon,\quad\varepsilon>0$ (16)
Here, $\phi$ serves as a content–target gate head, where the sigmoid function $\sigma(\cdot)$ determines the target signal strength. In contrast, $\omega_{\tau}$ is a strictly positive time–constant gate head that controls the rate of convergence and the steady-state amplitude. Conceptually, this parallels recurrent gating: $\phi$ regulates what content to emphasize, while $\omega_{\tau}$ governs how quickly and to what extent it is expressed.
Input Curation: We experimented with different strategies for constructing query–key inputs. Initially, we implemented full pairwise concatenation, where queries $Q\in\mathbb{R}^{B\times H\times T_{q}\times D}$ are combined with all keys $K\in\mathbb{R}^{B\times H\times T_{k}\times D}$ to form a joint tensor $U\in\mathbb{R}^{B\times H\times T_{q}\times T_{k}\times 2D}$. While this preserved complete feature information and enabled expressive, learnable similarity functions, it was memory-intensive, making it impractical for longer sequences. To mitigate this, we applied a sparse Top-K optimization: for each query, we compute pairwise scores $S=Q\cdot K^{\top}\in\mathbb{R}^{B\times H\times T_{q}\times T_{k}}$, select the Top-$K_{\text{eff}}=\min(K,T_{k})$ keys, and construct concatenated pairs $U_{\text{topk}}\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times 2D}$. This approach preserves the most relevant interactions while substantially reducing memory requirements in the concatenation and subsequent backbone processing stages, allowing the method to scale linearly with the sequence length in those components. However, the initial computation of $S$ remains quadratic (see Appendix C.3). Algorithm 2 outlines the steps required for input curation.
Time Vector: NAC builds on continuous-depth models as (hasani2022closed) that adapt their temporal dynamics to the task. It constructs an internal, normalized pseudo-time vector $t_{\text{pseudo}}$ using a sigmoidal transformation, $t_{\text{pseudo}}=\sigma(t_{a}\cdot t+t_{b})$, where $t_{a}$ and $t_{b}$ are learnable affine parameters and $\sigma$ is the sigmoid function. For time-varying datasets (e.g., irregularly sampled series), each time point $t$ is derived from the sample’s timestamp, while for tasks without meaningful timing, $t$ is set to $1$. The resulting $t_{\text{pseudo}}$ lies in $[0,1]$ and provides a smooth, bounded representation of time for modulating the network’s dynamics.
Attention logits and weights: Starting from Eqn. 3, consider the trajectory of a query–key pair with initial condition $a_{0}=0$:
$a_{t}=\frac{\phi}{\omega_{\tau}}\left(1-e^{-\omega_{\tau}t}\right),$ (17)
For finite $t$, the exponential factor $(1-e^{-\omega_{\tau}t})$ regulates the buildup of attention, giving $\omega_{\tau}$ a temporal gating role. Normalizing across all keys via softmax yields attention weights $\alpha_{t}=\text{softmax}(a_{t})$, defining a valid probability distribution where $\phi$ amplifies or suppresses content alignments, and $\omega_{\tau}$ shapes both the speed and saturation of these preferences.
As $t\to\infty$, the trajectory converges to the steady state
$a^{*}_{t}=\frac{\phi}{\omega_{\tau}}\approx\frac{q^{\top}k}{\sqrt{d_{k}}},$ (18)
which is analogous to scaled-dot attention under specific parameterization when the backbone $\mathcal{NN}_{\text{backbone}}$ is configured as a linear projection such that $\phi(\mathbf{u})=q^{\top}k$ and $\omega_{\tau}(u)=\sqrt{d_{k}}$ (e.g., by setting NCP weights to emulate bilinear forms and disabling nonlinearities). In general, the nonlinear backbone allows for more expressive similarities, with the approximation holding when trained to mimic dot products.
Attention output: Finally, the attention output is computed by integrating the attention weights with the value matrix:
$\text{NAC}(q,k,v)=\int_{T}\alpha_{t}v_{t}dt$ (19)
In practice, the integration is approximated using a Riemann-style approach, where the weighted elements are computed by multiplying each $v_{t}$ with its corresponding $\alpha_{t}$. These are then summed and multiplied by a fixed pseudo-time step $\delta_{t}$, chosen as a scalar (typically between 0.5–1.0) hyperparameter during layer initialization. This yields a continuous analogue of standard weighted sums, giving finer resolution of the attention trajectory without altering the underlying values. Sensitivity to attention output w.r.t $\delta_{t}$ is visualized in Appendix D.2.
2.2.1 Extension to Multi-head
To scale this mechanism to multi-head attention, we project the input sequence into $H$ independent subspaces (heads) of dimension $d_{\text{model}}/H$, yielding query, key, and value tensors $(q^{(h)},k^{(h)},v^{(h)})$ for $h\in\{1,\dots,H\}$. For each head, pairwise logits are computed according to Eqns. 2,3 or 18, followed by the softmax normalization to calculate attention weights. The resulting attention weights $\alpha^{(h)}_{t}$ are then used to integrate with the value vector $v^{(h)}$, producing head-specific attention outputs. Finally, these outputs are concatenated and linearly projected back into the model dimension. This formulation ensures that each head learns distinct dynamic compatibilities governed by its own parameterization of $\phi$ and $\omega_{\tau}$, while the aggregation across heads preserves the expressive capacity of the standard multi-head attention mechanism.
Figure 2: Illustration of the architecture of (a) Neuronal Attention Circuit mechanism ; (b) Multi-Head Extension
2.3 NAC as Universal Approximator
We now establish the universal approximation capability of NAC by extending the classical Universal Approximation Theorem (UAT) (nishijima2021universal) to the proposed mechanism. For brevity, we consider a network with a single NAC layer processing fixed-dimensional inputs, though the argument generalizes to sequences.
Theorem 2 (Universal Approximation by NAC).
Let $K\subset\mathbb{R}^{n}$ be a compact set and $f:K\to\mathbb{R}^{m}$ be a continuous function. For any $\epsilon>0$, there exists a neural network consisting of a single NAC layer, with sufficiently large model dimension $d_{\text{model}}$, number of heads $H$, sparsity $s$, and nonlinear activations, such that the network’s output $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfies
$\sup_{x\in K}\|f(x)-g(x)\|<\epsilon.$ (20)
The proof is provided in Appendix B.3.
3 Evaluation
We evaluate the proposed architecture against a range of baselines, including (DT & CT) RNN, (DT & CT) attention, and multiple NAC ablation configurations. Experiments are conducted across diverse domains, including irregular time-series modeling, lane keeping of autonomous vehicles, and Industry 4.0 prognostics. All results are obtained using 5-fold cross-validation, where models are trained using BPTT (see Appendix C.2) on each fold and evaluated across all folds. We report the mean ($\mu$) and standard deviation ($\sigma$) to capture variability and quantify uncertainty in the predictions. Table 1 provides results for all experiments, and the details of the baselines, ablation, environment utilized, the data curation and preprocessing, and neural network architectures for all experiments are provided in the Appendix D.3.
Table 1: Model Performance Across All Categories and Datasets
Model Irregular Time-Series Lane-Keeeping of AVs Industry 4.0
E-MNIST (↑)
PAR (↑) CarRacing (↑)
Udacity (↓)
PRONOSTIA (↓) XJTU-SY (↓) HUST (↓)
[…]
Note: (↑) higher is better; (↓) lower is better.
3.1 Irregular Time-series
We evaluate the proposed architecture on two irregular time-series datasets: (i) Event-based MNIST; and (ii) Person Activity Recognition (PAR).
Event-based MNIST: Event-based MNIST is the transformation of the widely recognized MNIST dataset with irregular sampling added originally proposed in (lechner2022mixed). The transformation was done in two steps: (i) flattening each 28×28 image into a time series of length 784, and (ii) encoding the binary time series into an event-based format by collapsing consecutive identical values (e.g., 1,1,1,1 → (1, t=4)). This representation requires models to handle temporal dependencies effectively. NAC-PW achieved first place with an accuracy of 96.64%, followed by NAC-Exact/05s/8k at 96.12%. GRU-ODE and ContiFormer ranked third with 96.04%.
Person Activity Recognition (PAR): We employed the Localized Person Activity dataset from UC Irvine (localization_data_for_person_activity_196). The dataset contains data from five participants, each equipped with inertial measurement sensors sampled every 211 ms. The goal of this experiment is to predict a person’s activity from a set of predefined actions, making it a classification task. All models performed well on this task, with NAC-PW achieving 89.15% accuracy and taking first place. NAC-Exact/05s/8k and GRU-ODE ranked second with 89.01% accuracy, while NAC-02s ranked third with 88.84% mean accuracy.
3.2 Lane-Keeping of Autonomous Vehicles
Lane keeping in autonomous vehicles (AVs) is a fundamental problem in robotics and AI. Early works (tiang2018lane; park2021convolutional) primarily emphasized accuracy, often relying on large models. More recent research (lechner2020neural; razzaq2023neural) has shifted toward designing compact architectures suitable for resource-constrained devices. The goal of this experiment is to create a long causal structure between the road’s horizon and the corresponding steering commands. To evaluate, we used two widely adopted simulation environments: (i) OpenAI CarRacing (brockman2016openai); and (ii) the Udacity Self-Driving Car Simulator (udacity). In OpenAI CarRacing, the task is to classify steering actions from a predefined action set. In contrast, the Udacity Simulator requires predicting a continuous trajectory of steering values. We implemented the AI models proposed by (razzaq2023neural), replacing the recurrent layer with NAC and its counterparts. All models achieved around 80% accuracy on average in the CarRacing benchmark. Notably, NAC-PW performed the best, reaching the highest accuracy of 80.72%, followed by NAC-Steady, ranked second with 80.62%. LSTM and GRU took third position, achieving 80.60% on average. In the Udacity benchmark, NAC-32k performed the best, achieving the lowest MSE of 0.0170. NAC-Exact followed with 0.0173, and ContiFormer ranked third with 0.0174. To visualize saliency maps for these experiments, refer to Appendix D.3.3. Experimental videos are available for the OpenAI CarRacing [click here] and for the Udacity Simulator [click here].
Figure 3: Degradation Estimation Results.
3.3 Industry 4.0
Industry 4.0 has transformed manufacturing, making prognostic health management (PHM) systems essential. A key PHM task is estimating the remaining useful life (RUL) of components, particularly rolling element bearings (REB), which account for 40–50% of machine failures (ding2021remaining; zhuang2021temporal). The objective is to learn degradation features from one operating condition of a dataset and generalize to unseen conditions within the same dataset. Furthermore, the model should provide accurate RUL estimation on entirely different datasets, while maintaining a compact architecture suitable for resource-constrained devices, thereby supporting localized safety.
We utilized three benchmark datasets: (i) PRONOSTIA (nectoux2012pronostia), (ii) XJTU-SY (wang2018xjtu), and (iii) HUST (thuan2023hust). Training is performed on PRONOSTIA, while XJTU-SY and HUST are used to assess cross-validation. We used the Score metric (nectoux2012pronostia) to assess the performance. We evaluate generalization using Bearing 1 from the first operating condition of each dataset. Figure 3 visualizes the expected degradation alongside the outputs of NAC. On the PRONOSTIA dataset, ContiFormer performed the best with the lowest score of 27.82. NAC-Exact/05s/8k and NAC-PW achieved nearly identical scores, obtaining 37.75 and 37.50 on average, respectively. On the XJTU-SY dataset, NAC-Exact/05s/8k has the lowest score of 19.87. NAC-FC ranked second with a score of 24.89, followed by NAC-PW in third place with an average score of 28.01. A similar trend was observed on the HUST dataset, where NAC-Exact/05s/8k achieved first place with a score of 27.82, NAC-PW ranked second with 30.14, and NAC-FC ranked third with 35.35. These results demonstrated the strong cross-validation capability of NAC.
Table 2: Run-Time and Memory Benchmark Results
Model
Run-Time
Throughput
(seq/s)
Peak Memory
RNN $1.8392^{\scriptsize\pm 0.1933}$ 0.544 0.29
CT-RNN $7.1097^{\scriptsize\pm 0.3048}$ 0.141 0.67
LSTM $2.6241^{\scriptsize\pm 0.2906}$ 0.381 0.42
PhasedLSTM $4.9812^{\scriptsize\pm 0.272}$ 0.201 0.80
GRU $3.216^{\scriptsize\pm 0.2566}$ 0.311 0.54
GRU-ODE $12.2498^{\scriptsize\pm 0.0525}$ 0.082 0.64
mmRNN $7.5852^{\scriptsize\pm 0.2785}$ 0.132 0.96
LTC $14.643^{\scriptsize\pm 0.2445}$ 0.068 0.99
CfC $6.0988^{\scriptsize\pm 0.2135}$ 0.164 0.76
Attention $0.0016^{\scriptsize\pm 0.0001}$ 625.00 16.86
MHA $0.0041^{\scriptsize\pm 0.0001}$ 243.90 69.05
mTAN $0.0272^{\scriptsize\pm 0.0054}$ 36.76 790.16
ODEFormer $0.0317^{\scriptsize\pm 0.0016}$ 31.55 67.71
CTA $8.5275^{\scriptsize\pm 0.2355}$ 0.117 1.43
ContiFormer $0.066^{\scriptsize\pm 0.0075}$ 15.15 67.71
NAC-2k $7.3071^{\scriptsize\pm 0.1547}$ 0.137 44.75
NAC-32k $7.2313^{\scriptsize\pm 0.219}$ 0.138 549.86
NAC-PW $8.5649^{\scriptsize\pm 0.0203}$ 0.117 5042.09
NAC-FC $0.0195^{\scriptsize\pm 0.0002}$ 51.28 29.92
NAC-02s $7.252^{\scriptsize\pm 0.2018}$ 0.138 151.54
NAC-09s $7.222^{\scriptsize\pm 0.176}$ 0.139 150.85
NAC-Exact/05s/8k $7.4101^{\scriptsize\pm 0.1586}$ 0.135 151.50
NAC-Euler $7.3367^{\scriptsize\pm 0.1719}$ 0.136 152.22
NAC-Steady $7.2942^{\scriptsize\pm 0.1451}$ 0.137 150.86
3.4 Run Time and Memory Experiments
We evaluate computational requirements on fixed-length sequences of 1024 steps, 64-dimensional features, 4 heads, and a Batch size of 1. Each model is run for ten forward passes on Google Colab T4-GPU, and we report the mean runtime with standard deviation, throughput, and peak memory usage. NAC occupies an intermediate position in runtime relative to several CT-RNN models, including GRU-ODE, CfC, and LTC. In terms of memory consumption, NAC uses significantly less memory than mTAN, with NAC-2k being the least memory-consuming among the CT-Attention models. Reducing NAC sparsity from 90% to 20% has minimal effect on memory, decreasing usage slightly from 151.54 MB to 150.85 MB. In constrast, decreasing the Top-K selection from PW to $k=2$ drastically reduces memory consumption from 5042 MB to 44.75 MB, demonstrating the flexibility of NAC.
Interpreting the Results: We observe that increasing sparsity improves system robustness and leads to higher overall accuracy. Similarly, increasing Top-$K$ interactions improves accuracy, though the benefits diminish with growing memory consumption. We recommend a Top-$K$ of 8 with a 50% threshold as a robust default for this trade-off. In general, for tasks requiring adaptive continuous-time dynamics due to evolving temporal patterns, use the transient mode (Exact or Euler). For tasks where temporal dynamics are negligible, use the Steady mode for fast inference.
4 Discussions
This paper is part of ongoing research on biologically plausible attention mechanisms and represents a pioneering step, with limitations to be addressed in future work.
Architectural improvement: Currently, NAC uses pre-determined wiring (AutoNCP) requiring three inputs: number of units (sensory + interneuron + motor), output motor neurons, and sparsity, with typically 60% of units assigned to sensory neurons. To integrate with the attention mechanism while preserving wiring, sensory units for $\mathcal{NN_{\text{sensory}}}$ are set as $\text{units}_{\text{sensory}}=\left\lceil\frac{d_{\text{model}}-0.5}{0.6}\right\rceil$ and backbone units as $\text{units}_{\textit{backbone}}=d_{\text{model}}+\left\lfloor\frac{d_{\text{model}}}{0.6}\right\rfloor$, where $\lceil\cdot\rceil$ and $\lfloor\cdot\rfloor$ denote the ceiling and floor functions, respectively. This results in a larger overall architectural size and increased runtime. Future work will support user-defined NCPs configurations or randomized wiring to enable more efficient architectures.
Learnable sparse Top-K selection: Sparse Top-K attention can miss important context, is sensitive to $k$, and may be harder to optimize. A further limitation is that it still computes the full $QK^{\top}$ matrix, which can dominate the cost for very long sequences. Future work includes adaptive or learnable Top-K selection, improved key scoring, and hardware-aware optimization to strengthen accuracy and robustness.
5 Conclusion
In this paper, we introduce the Neuronal Attention Circuit (NAC), a biologically inspired attention mechanism that reformulates attention logits as the solution to a first-order ODE modulated by nonlinear, interlinked gates derived from repurposing C.elegans nematode NCPs. NAC bridges discrete attention with continuous-time dynamics, enabling adaptive temporal processing without the limitations inherent to traditional scaled dot-product attention. Based on the solution to ODE, we introduce three computational modes: (i) Euler, based on explicit Euler integration; (ii) Exact, providing closed-form solutions; and (iii) Steady, approximating equilibrium states. In addition, a sparse Top-K pairwise concatenation scheme is introduced to mitigate the memory intensity. Theoretically, we establish NAC’s log-state stability, exponential error bounds, and universal approximation, thereby providing rigorous guarantees of convergence and expressiveness. Empirical evaluations demonstrate that NAC achieves state-of-the-art performance across diverse tasks, including irregularly sampled time-series benchmarks, autonomous vehicle lane-keeping, and industrial prognostics. Moreover, NAC occupies an intermediate position between CT-RNNs and CT-Attention, offering robust temporal modeling while requiring less runtime than CT-RNNs and less memory than CT-Attention models.
Reproducibility Statement
The code for reproducibility is available at https://github.com/itxwaleedrazzaq/neuronal_attention_circuit
Impact Statement
The work addresses the growing field of continuous-time attention and pioneers a biologically plausible mechanism. It encourages research into sparse, adaptive networks that resemble natural wiring. From a societal perspective, it supports more robust AI in resource-limited settings, but it also raises ethical concerns when applied to areas such as surveillance or autonomous systems.
Appendix
Appendix A Preliminaries
A.1 Attention Mechanism
Attention mechanisms have become a cornerstone in modern neural architectures, enabling models to dynamically focus on relevant parts of the input. The concept was first introduced in the context of neural machine translation, where it allowed the decoder to weight encoder outputs according to their importance for generating each target token. Formally, given a query vector $q\in\mathbb{R}^{d}$, key vectors $K=[k_{1},k_{2},\dots,k_{n}]\in\mathbb{R}^{n\times d}$, and value vectors $V=[v_{1},v_{2},\dots,v_{n}]\in\mathbb{R}^{n\times d}$, the attention mechanism can be expressed in two steps:
  1. 1.
    Compute the scaled dot attention logits:
    $a_{i}=\frac{q^{T}k_{i}}{\sqrt{d}}$ (21)
  2. 2.
    Normalize the logits to get attention weights and compute the output:
    $\alpha_{i}=\text{softmax}(a_{i})$ $=\frac{e^{a_{i}}}{\sum_{j=1}^{n}e^{a_{j}}}$ (22)
    $\text{Attention}(q,k,v)$ $=\sum_{i=1}^{n}\alpha_{i}v_{i}$ (23)
Here, $a_{i}$ is the raw attention logit between the query and each key, and the scaling factor $\sqrt{d}$ prevents large dot products from destabilizing the softmax (vaswani2017attention).
A.2 Neuronal Circuit Policies (NCPs)
NCPs represent a biologically inspired framework for developing interpretable neural control agents by adapting the tap-withdrawal circuit found in the nematode C. elegans (lechner2018neuronal). Unlike traditional spiking neural networks, the majority of neurons in this circuit exhibit electronic dynamics, characterized by the passive flow of electrical charges, resulting in graded potentials. NCPs are structured as a four-layer hierarchical architecture comprising sensory neurons ($N_{s}$), interneurons ($N_{i}$), command neurons ($N_{c}$), and motor neurons ($N_{m}$). The $N_{s}$ perceive and respond to external stimulus inputs and are responsible for the initial signal transduction. Each $N_{s}$ consists of subneurons $S_{p}$ and $S_{n}$ and a system variable $x$. The activation of $S_{p}$ and $S_{n}$ depends upon the sign of $x$: $S_{p}$ becomes activated for $x>0$, whereas $S_{n}$ becomes activated for $x<0$. The variable $x$ is mapped to the membrane potential range of $70\,\text{mV},-20\,\text{mV}]$, which is consistent with the biophysical behavior of nerve cells, which typically exhibit a resting potential at $70\,\text{mV}$ and an activation potential near $20\,\text{mV}$. Similarly, each $N_{m}$ is composed of two subneurons, $M_{p}$ and $M_{n}$, and is driven by a controllable variable $y$, which also maps to a biologically plausible range $70\,\text{mV},-20\,\text{mV}]$. The connections in the NCP architecture are designed to reflect the biological sparsity and abstraction of neural circuits. Specifically, connections from $N_{s}$ to $N_{i}$ are feedforward, while those between $N_{c}$ and $N_{m}$ are highly recurrent (lechner2018neuronal). Figure 1(a) illustrates the connectome of NCPs.
Appendix B Proofs
In this section, we provide all the proofs.
B.1 Deriving Closed-form (Exact) Solution
Although $\phi$ and $\omega_{\tau}$ are nonlinear functions of the input $\mathbf{u}=[\mathbf{q};\mathbf{k}]$, we derive closed-form solution by treating them as locally constant over the pseudo-time integration interval for each query–key pair based on frozen-coefficient approximation (john1952integration). This is accurate whenever the interval is short or when input variations are slow compared with the relaxation rate $\omega_{\tau}$. Under approximation assumption, rewrite Eqn. 1 as
$\frac{da_{t}}{dt}+\omega_{\tau}a_{t}=\phi.$ (24)
This is now a linear first-order ODE. The integrating factor is
$\mu=e^{\Big(\int\omega_{\tau}\,dt\Big)}=e^{\omega_{\tau}t}.$ (25)
Multiply both sides by $\mu(t)$:
$e^{\omega_{\tau}t}\frac{da_{t}}{dt}+\omega_{\tau}e^{\omega_{\tau}t}a_{t}=\phi e^{\omega_{\tau}t}.$ (26)
Recognize the left-hand side as the derivative of $e^{\omega_{\tau}t}a_{t}$:
$\frac{d}{dt}\big(e^{\omega_{\tau}t}a_{t}\big)=\phi e^{\omega_{\tau}t}.$ (27)
Integrate from 0 to $t$:
$e^{\omega_{\tau}t}a_{t}-e^{0}a_{0}=\phi\int_{0}^{t}e^{\omega_{\tau}s}\,ds.$ (28)
Compute the integral (since $\omega_{\tau}\neq 0$):
$\int_{0}^{t}e^{\omega_{\tau}s}\,ds=\frac{1}{\omega_{\tau}}\big(e^{\omega_{\tau}t}-1\big).$ (29)
Substitute back:
$e^{\omega_{\tau}t}a_{t}-a_{0}=\phi\cdot\frac{e^{\omega_{\tau}t}-1}{\omega_{\tau}}.$ (30)
Rearrange:
$e^{\omega_{\tau}t}a_{t}=a_{0}+\frac{\phi}{\omega_{\tau}}\big(e^{\omega_{\tau}t}-1\big).$ (31)
Divide both sides by $e^{\omega_{\tau}t}$:
$a_{t}=a_{0}e^{-\omega_{\tau}t}+\frac{\phi}{\omega_{\tau}}\big(1-e^{-\omega_{\tau}t}\big).$ (32)
Set $a^{*}:=\dfrac{\phi}{\omega_{\tau}}$. Then $a_{t}=a^{*}+(a_{0}-a^{*})e^{-\omega_{\tau}t}$, proved.
B.2 Proof of Theorem 1
We divide the proof into two parts: (i) the single-connection case $M=1$, and (ii) the general multi-connection case $M>1$. The main technique is to evaluate the ODE at boundary values of the proposed invariant interval and show that the derivative points inward, ensuring that trajectories cannot escape.
Case 1: Single connection ($M=1$).
The ODE reduces to
$\frac{da}{dt}=-\omega_{\tau}a+\phi=-\omega_{\tau}(a-A),\quad A=\frac{\phi}{\omega_{\tau}}.$ (33)
Here $A$ is the unique equilibrium. We now check both bounds.
- Upper bound: Let $M=\max(0,A)$. At $a=M$,
$\frac{da}{dt}\Big|_{a=M}=-\omega_{\tau}(M-A).$ (34)
If $A\geq 0$, then $M=A$ and $\frac{da}{dt}=0$. If $A<0$, then $M=0$, so $\frac{da}{dt}=-\omega_{\tau}(0-A)=\omega_{\tau}A\leq 0$ since $A<0$. In both cases, $\frac{da}{dt}\leq 0$. Thus trajectories cannot cross above $M$.
- Lower bound: Let $m=\min(0,A)$. At $a=m$,
$\frac{da}{dt}\Big|_{a=m}=-\omega_{\tau}(m-A).$ (35)
If $A\leq 0$, then $m=A$ and $\frac{da}{dt}=0$. If $A>0$, then $m=0$, so $\frac{da}{dt}=-\omega_{\tau}(0-A)=\omega_{\tau}A\geq 0$. In both cases, $\frac{da}{dt}\geq 0$. Thus trajectories cannot cross below $m$. Therefore, the interval $[m,M]$ is forward-invariant.
To see this explicitly under Euler discretization with step size $\Delta t>0$,
$a(t+\Delta t)=a_{t}+\Delta t\cdot\frac{da}{dt}.$ (36)
At $a=M$, $\frac{da}{dt}\leq 0\implies a(t+\Delta t)\leq M$. At $a=m$, $\frac{da}{dt}\geq 0\implies a(t+\Delta t)\geq m$. By induction over steps, $a_{t}\in[m,M]$ for all $t\in[0,T]$.
Case 2: Multiple connections ($M>1$).
The ODE is
$\frac{da}{dt}=-\Big(\sum_{j=1}^{M}f_{j}\Big)a+\sum_{j=1}^{M}f_{j}A_{j},$ (37)
with per-connection equilibria $A_{j}=\phi_{j}/f_{j}$. The effective equilibrium is
$A=\frac{\sum_{j=1}^{M}f_{j}A_{j}}{\sum_{j=1}^{M}f_{j}}.$ (38)
Since the weights $\frac{f_{j}}{\sum f_{j}}$ are positive and sum to 1, $A$ is a convex combination of $\{A_{j}\}$. Therefore,
$A\in[A^{\min},A^{\max}].$ (39)
- Upper bound: Let $M=\max(0,A^{\max})$. Then
$\frac{da}{dt}\Big|_{a=M}=\sum_{j=1}^{M}f_{j}(A_{j}-M).$ (40)
Since $A_{j}\leq A^{\max}\leq M$, each term $(A_{j}-M)\leq 0$, and thus $\sum f_{j}(A_{j}-M)\leq 0$. Hence $\frac{da}{dt}\leq 0$, proving trajectories cannot exceed $M$.
- Lower bound: Let $m=\min(0,A^{\min})$. Then
$\frac{da}{dt}\Big|_{a=m}=\sum_{j=1}^{M}f_{j}(A_{j}-m).$ (41)
Since $A_{j}\geq A^{\min}\geq m$, each $(A_{j}-m)\geq 0$, so $\sum f_{j}(A_{j}-m)\geq 0$. Hence $\frac{da}{dt}\geq 0$, proving trajectories cannot fall below $m$. Thus, the interval $[m,M]$ is forward-invariant.
Remark 1.
This result guarantees that the continuous-time attention state converges within a well-defined interval dictated by the per-connection equilibria. In particular, for the single-connection case ($M=1$), the state trajectory converges monotonically toward the closed-form equilibrium solution (Eqn. 18) without overshoot.
B.3 Proof for Theorem 2
The proof proceeds constructively by showing that the NAC layer can emulate a single-hidden-layer feedforward neural network with nonlinear activations, which is a universal approximator under the Universal Approximation Theorem (UAT). We assume self-attention on a single-token input $x\in\mathbb{R}^{n}$ (setting sequence length $T=1$) and focus on the steady mode for simplicity. Without loss of generality, set $d_{\text{model}}=n+m$ or adjust as needed for dimensionality. Constructively, set NCP sparsity $s=0$ for full connectivity, ensuring the backbone $\mathcal{NN}_{\text{backbone}}$ approximates any $\tilde{\phi}:\mathbb{R}^{2d}\to[0,1]$ with error $\delta$ via stacked layers. For multi-head, scale $H$ proportionally to target complexity, with output projection $W_{o}$ aggregating as in classical UAT proofs (stinchcomb1989multilayered).
Input Projections: The input $x$ is projected via NCP-based sensory projections to obtain query $q=q_{\text{proj}}(x)$, key $k=k_{\text{proj}}(x)$, and value $v=v_{\text{proj}}(x)$, each in $\mathbb{R}^{d_{\text{model}}}$. For emulation, set $q_{\text{proj}}=k_{\text{proj}}=I_{n}$ (identity on $\mathbb{R}^{n}$) and adjust $W_{v}$ later, so $q=k=x$ and $v$ is configurable. These are affine transformations and do not limit expressivity.
Head Splitting and Sparse Top-k Pairwise Computation: Split into $H$ heads, yielding $q^{(h)},k^{(h)}\in\mathbb{R}^{d}$ per head $h$, where $d=d_{\text{model}}/H$. For $T=1$, compute sparse top-k pairs, but since $T=1$, $K_{\text{eff}}=1$, yielding concatenated pair $u^{(h)}=[q^{(h)};k^{(h)}]\in\mathbb{R}^{2d}$. Since $q^{(h)}=k^{(h)}$, this is $x^{(h)};x^{(h)}]$, but the NCP processes it generally.
Computation of $\phi^{(h)}$ and $\omega_{\tau}^{(h)}$: The scalar $\phi^{(h)}$ is computed via the NCP-based inter-to-motor projection on the pair:
$\phi^{(h)}=\sigma(\mathcal{NN}_{\text{backbone}}(u^{(h)}))$ (42)
where $\sigma(z)=(1+e^{-z})^{-1}$ is the sigmoid. This NCP, with sufficiently large units and low sparsity, approximates any continuous scalar function $\tilde{\phi}:\mathbb{R}^{2d}\to[0,1]$ to arbitrary precision on compact sets (by the UAT for multi-layer networks (stinchcomb1989multilayered)). Similarly, $\omega_{\tau}^{(h)}$ is computed via:
$\omega_{\tau}^{(h)}=\mathrm{softplus}(\mathcal{NN}_{\text{backbone}}(u^{(h)}))+\varepsilon,\quad\varepsilon>0$ (43)
By setting weights to make $\omega_{\tau}^{(h)}\equiv 1$ (constant), the steady-mode logit simplifies to $a^{(h)}=\phi^{(h)}/\omega_{\tau}^{(h)}=\phi^{(h)}$. Thus, $a^{(h)}\approx\sigma\big(w^{(h)}x+b^{(h)}\big)$ for chosen weights $w^{(h)},b^{(h)}$, emulating a sigmoid hidden unit.
Attention Weights and output: For $T=1$, the softmax over one “key” yields $\alpha^{(h)}=\exp(a^{(h)})/\exp(a^{(h)})=1$. The head output is $y^{(h)}=\int_{T}\alpha^{(h)}v^{(h)}dt$. Set $v_{\text{proj}}$ such that $v^{(h)}=1$ (scalar), yielding $y^{(h)}\approx\sigma\big(w^{(h)}x+b^{(h)}\big)$. For vector-valued $v^{(h)}$, more complex combinations are possible, but scalars suffice here.
Output Projection: Concatenate head outputs: $Y=[y^{(1)};y^{(2)};\dots;y^{(H)}]\in\mathbb{R}^{H}$. Apply the final dense layer:
$g(x)=(Y\cdot W_{o})+b_{o}\in\mathbb{R}^{m}.$ (44)
With $y^{(h)}\approx\sigma\big(w^{(h)}x+b^{(h)}\big)$, this matches a single-hidden-layer network with $H$ units. By the UAT, for large $H$, such networks approximate any continuous $f$ on compact $K$ to accuracy $\epsilon$, by choosing appropriate $w^{(h)},b^{(h)},W_{o},b_{o}$.
Appendix C Training, Gradients and Complexity
C.1 Gradient Characterization
We analyze the sensitivity of the dynamics with respect to the underlying learnable parameters. Specifically, we compute closed-form derivatives of both the steady state and the full trajectory $a_{t}$ with respect to the parameters $\phi$ and $\omega_{\tau}$. These expressions illuminate how gradients flow through the system, and provide guidance for selecting parameterizations that avoid vanishing or exploding gradients.
C.1.1 Trajectory sensitivities for Closed-form formulation
The trajectory is given by
$a_{t}=a^{*}+(a_{0}-a^{*})e^{-\omega_{\tau}t},$ (45)
which depends on $(\phi,\omega_{\tau})$ both through the equilibrium $a^{*}$ and the exponential term.
Derivative with respect to $\phi$: We obtain
$\frac{\partial a_{t}}{\partial\phi}=\frac{1-e^{-\omega_{\tau}t}}{\omega_{\tau}}$ (46)
Interpretation: For large $\omega_{\tau}$, the gradient with respect to $\phi$ saturates quickly but shrinks to scale $\mathcal{O}(1/\omega_{\tau})$, potentially slowing learning of $\phi$. Conversely, very small $\omega_{\tau}$ leads to large steady-state gradients, which may destabilize optimization.
Derivative with respect to $\omega_{\tau}$: Here, both the equilibrium and the decay rate depend on $\omega_{\tau}$, yielding
$\frac{\partial a_{t}}{\partial\omega_{\tau}}=-\frac{\phi}{\omega_{\tau}^{2}}\big(1-e^{-\omega_{\tau}t}\big)-(a_{0}-a^{*})\,t\,e^{-\omega_{\tau}t}.$ (47)
Interpretation: The gradient with respect to $\omega_{\tau}$ contains a transient term proportional to $te^{-\omega_{\tau}t}$, which dominates at intermediate times, and a steady-state contribution proportional to $\phi/\omega_{\tau}^{2}$, which persists asymptotically. Thus, sensitivity to $\omega_{\tau}$ is time-dependent, peaking before vanishing exponentially in the transient component.
C.2 Gradient-Based Training
Like Neural ODEs (chen2018neural) and CT-RNNs (rubanova2019latent), NAC produces differentiable computational graphs and can be trained using gradient-based optimization, such as the adjoint sensitivity method (cao2003adjoint) or backpropagation through time (BPTT) (lecun1988theoretical). In this work, we use BPTT exclusively, as the adjoint sensitivity method can introduce numerical errors (zhuang2020adaptive).
C.3 Efficiency and Complexity
Table 3 summarizes the computational complexity of different sequence models. For sequence prediction over length $n$ with hidden dimension $k$, RNNs scale linearly, $\mathcal{O}(nk)$, while Attention and NAC scale quadratically, $\mathcal{O}(n^{2}k)$. ODE-based models, such as LNNs, incur an additional multiplicative factor $S$ for the number of solver steps. For single-time-step prediction, RNNs and LSTMs require $\mathcal{O}(k)$, whereas Attention and NAC require $\mathcal{O}(nk)$ when recomputing attention over the full sequence.
Table 3: Sequence and time-step prediction complexity. $n$ is the sequence length and $k$ is the hidden/model dimension.
Model Sequence Time-step
RNN $\mathcal{O}(nk)$ $\mathcal{O}(k)$
Attention $\mathcal{O}(n^{2}k)$ $\mathcal{O}(nk)$
LNN (ODEsolve) $\mathcal{O}(nk\cdot S)$ $\mathcal{O}(k\cdot S)$
NAC-Exact $\mathcal{O}(n^{2}k)$ $\mathcal{O}(nk)$
NAC-Euler $\mathcal{O}(n^{2}k)$ $\mathcal{O}(nk)$
Appendix D Evaluation
D.1 Related Works
The brief description for related works is divided into four subcategories.
DT-RNNs: RNN (rumelhart1985learning) captures sequential dependencies in time-series data by updating a hidden state from the current observation and the previous state. LSTM (hochreiter1997long) extends RNNs with input, output, and forget gates, allowing the network to maintain and update long-term memory, which improves modeling of long-term dependencies in time-series sequences. GRU (cho2014learning) simplifies the LSTM architecture by combining the forget and input gates into a single update gate, allowing efficient modeling of long-term dependencies in time-series sequences.
CT-RNNs: CT-RNN (rubanova2019latent) model temporal dynamics using differential equations, enabling hidden states to evolve continuously over time in response to inputs, which is particularly useful for irregularly sampled time-series data. PhasedLSTM (neil2016phased) introduces a time gate that updates hidden states according to a rhythmic schedule, enabling efficient modeling of asynchronous or irregularly sampled time-series. GRU-ODE (de2019gru) extends the GRU to continuous time, evolving hidden states via ODEs to handle sequences with non-uniform time intervals. mmRNN (lechner2022mixed) combines short-term and long-term memory units to capture both fast-changing and slowly evolving patterns in sequential data. LTC (hasani2021liquid) use neurons with learnable, input-dependent time constants to adapt the speed of dynamics and capture complex temporal patterns in continuous-time data. CfC (hasani2022closed) approximate LTC dynamics analytically, providing efficient continuous-time modeling without relying on numerical ODE solvers.
DT-Attentions: Attention (vaswani2017attention) computes attention weights by measuring similarity between queries and keys, scaling the results, and applying softmax to weigh time-step contributions. Multi-Head Attention (vaswani2017attention) applies multiple parallel scaled dot-product attention mechanisms, capturing different types of temporal dependencies simultaneously for complex time-series modeling.
CT-Attentions: mTAN (shukla2021multi) learns continuous-time embeddings and uses time-based attention to interpolate irregular observations into a fixed-length representation for downstream encoder-decoder modeling. CTA (chien2021continuous) generalizes discrete-time attention to continuous-time by representing hidden states, context vectors, and attention scores as functions whose dynamics are modeled via neural networks and integrated using ODE solvers for irregular sequences. ODEFormer(d2023odeformer) trains a sequence-to-sequence transformer on synthetic trajectories to directly output a symbolic ODE system from noisy, irregular time-series data. ContiFormer (chen2023contiformer) builds a continuous-time Transformer by pairing ODE-defined latent trajectories with a time-aware attention mechanism to model dynamic relationships in irregular time-series data.
D.2 Ablations Details
The brief descriptions of variants and ablation are also divided into four subcategories:
Top-K Ablations: NAC-2k uses Top-K=2 to compute the logits and NAC-32k uses Top-K=32. All variants use the exact computation mode with 50% sparsity.
Sparsity Ablations: NAC-02s uses 20% sparsity to compute the logits and NAC-09s uses 90%. NAC-PW employs full pairwise (non-sparse) concatenation for input curation. NAC-FC replaces the sparse NCP gating mechanism with a simple fully connected layer. All variants use the exact computation mode with Top-K=8.
Modes variants: NAC-Euler computes attention logits using the explicit Euler integration method. NAC-Steady derives attention logits from the steady-state solution of the exact formulation. NAC-Exact/05s/8k computes attention logits using the closed-form exact solution. It also overlaps with other ablations, so we combined it into a single one. All modes use Top-K=8, 50% sparsity and $\delta_{t}$=1.0. The sensitivity of NAC to $\delta_{t}$ is visualized in Figure 4
Figure 4: Effect of $\delta_{t}$ on output of NAC.
D.3 Experimental Details
D.3.1 Event-based MNIST
Dataset Explanation and Curation: The MNIST dataset, introduced by (deng2012mnist), is a widely used benchmark for computer vision and image classification tasks. It consists of 70,000 grayscale images of handwritten digits (0–9), each of size $28\times 28$ pixels, split into 60,000 training and 10,000 testing samples.
Preprocessing: We follow the preprocessing pipeline described in (lechner2022mixed), which proceeds as follows. First, a threshold is applied to convert the 8-bit pixel values into binary values, with 128 as the threshold on a scale from 0 (minimum intensity) to 255 (maximum intensity). Second, each $28\times 28$ image is reshaped into a one-dimensional time series of length 784. Third, the binary time series is encoded in an event-based format, eliminating consecutive occurrences of the same value; for example, the sequence $[1,1,1,1]$ is transformed into $(1,t=4)$. This encoding introduces a temporal dimension and compresses the sequences from 784 to an average of 53 time steps. Finally, to facilitate efficient batching and training, each sequence is padded to a fixed length of 256, and the time dimension is normalized such that each symbol corresponds to one unit of time. The resulting dataset defines a per-sequence classification problem on irregularly sampled time series.
Neural Network Architecture: We develop an end-to-end hybrid neural network by combining compact convolutional layers with NAC or counterparts baselines for fair comparison. Detailed hyperparameters and architectural specifications are provided in Table 4.
D.3.2 Person Activity Recognition (PAR)
Dataset Explanation and Curation: We used the Localized Person Activity Recognition dataset provided by UC Irvine (localization_data_for_person_activity_196). The dataset comprises 25 recordings of human participants performing different physical activities. The eleven possible activities are “walking,” “falling,” “lying down,” “lying,” “sitting down,” “sitting,” “standing up from lying,” “on all fours,” “sitting on the ground,” “standing up from sitting,” and “standing up from sitting on the ground.” The objective of this experiment is to recognize the participant’s activity from inertial sensors, formulating the task as a per-time-step classification problem. The input data consist of sensor readings from four inertial measurement units placed on participants’ arms and feet. While the sensors are sampled at a fixed interval of 211 ms, recordings exhibit different phase shifts and are thus treated as irregularly sampled time series.
Preprocessing: We first separated each participant’s recordings based on sequence identity and calculated elapsed time in seconds using the sampling period. To mitigate class imbalance, we removed excess samples from overrepresented classes to match the size of the smallest class. Subsequently, the data were normalized using a standard scaler. Finally, the dataset was split into a 90:10 ratio for training and testing.
Neural Network Architecture: Following the approach in Section D.3.1, we developed an end-to-end hybrid neural network combining convolutional heads with NAC or other baselines. Hyperparameter details are summarized in Table 4.
D.3.3 Autonomous Vehicle
Dataset Explanation and Curation: We followed the data collection methodology described in (razzaq2023neural). For OpenAI-CarRacing, a PPO-trained agent (5M timesteps) was used to record 20 episodes, yielding approximately 48,174 RGB images of size $92\times 92\times 3$ with corresponding action labels across five discrete actions (no-act, move left, forward, move right, stop). The dataset was split with 10% reserved for testing and the remaining 90% for training. For the Udacity simulator, we manually controlled the vehicle for 50 minutes, producing 15647 RGB images of size $320\times 160\times 3$, captured from three camera streams (left, center, right) along with their corresponding continuous steering values. This dataset was split into 20% testing and 80% training.
Preprocessing: No preprocessing was applied to the OpenAI-CarRacing dataset. For the Udacity simulator, we followed the preprocessing steps in (shibuya_car_behavioral_cloning). Each image was first cropped to remove irrelevant regions and resized to $66\times 120\times 3$. Images were then converted from RGB to YUV color space to match the network input. To improve robustness, data augmentation techniques, including random flips, translations, shadow overlays, and brightness variations, were applied to simulate lateral shifts and diverse lighting conditions.
Neural Network Architecture: For OpenAI-CarRacing, we modified the neural network architecture proposed in (razzaq2023neural), which combines compact CNN layers for spatial feature extraction with LNNs to capture temporal dynamics. In our implementation, the LNN layers were replaced with NAC and its comparable alternatives for fair evaluation. Full hyperparameter configurations are provided in Table 4. For the Udacity simulator, we modified the network proposed in (bojarski2016end) by replacing three latent MLP layers with NAC and its counterparts. Full hyperparameters for this configuration are summarized in Table 4.
Saliency Maps: A saliency map visualizes the regions of the input that a model attends to when making decisions. Figure 5 shows the saliency maps for the OpenAI CarRacing environment. We observe that only NAC (Steady, Euler, and Steady) maintains focus on the road’s horizon, while other models either focus on the sides or remain largely unresponsive to the task. Figure 6 also presents the saliency maps for the Udacity Simulator. In this case, NAC-Exact, CTA produces the most accurate visual maps, maintaining attention on the road’s horizon, followed by ContiFormer and mTAN, which achieve comparable performance. Attention and PhasedLSTM also generate reasonable saliency maps, although their focus is more dispersed across the scene. In contrast, other models either fail to identify relevant regions, producing blurry maps, or focus solely on one side of the road. These results demonstrate the NAC’s ability to understand the underlying task.
Figure 5: Saliency maps for OpenAI CarRacing
Figure 6: Saliency maps for Udacity Simulator
Table 4: Summary of Key Hyperparameters of All Experiments
Param. MNIST PAR CarRacing Udacity RUL EST.
Conv layers 1D(64@5) 1D(64@5, 64@3) 3×TD-2D(10–30@3–5) 2D(24–64@5–3, ELU) 1D(32@3, 16@2)
NAC 64-d, 8h 32-d, 4h 64-d, 16h 100-d, 16h 16-d, 8h
Dense 32–10(SM) 32–11(SM) 64–5(SM) 64–1(Lin) 1(Lin)
Dropout 0.2 0.5
Opt. AdamW AdamW Adam AdamW AdamW
LR 0.001 0.001 0.0001 0.001 0.001
Loss SCE SCE SCE MSE MSE
Metric Acc Acc Acc MAE Score
Batch 32 20 32 32
Epochs 150 500 100 10 150
Note: SCE = Sparse Categorical Crossentropy; Acc = Accuracy; MAE = Mean Absolute Error; MSE = Mean Squared Error; SM = softmax; Lin = Linear; TD = TimeDistributed; Conv1D/2D = Conv1D/2D; $d$ = model dimension; $h$ = attention heads.
Baselines Hyperparameters Clarification: All (CT & DT) RNNs use the same number of hidden units as NAC’s $d_{\text{model}}$, and all (DT & CT) Attention use the same $d_{\text{model}}$ and heads as NAC. The other layers, including 1D/2D, Dense, and the remaining hyperparameters, are the same during our tests.
D.3.4 Industry 4.0
Algorithm 3 Time‒frequency Representation Algorithm
0:  windowed signal $I_{w}$, critical frequency $f_{c}$, operating frequency $f_{o}$, sampling period $T_{sampling}$, windowed physical constraints ($t_{w},T_{w}$)
  $a_{min}=\frac{f_{c}}{f_{max}\cdot T_{sampling}},\quad a_{{max}}=\frac{f_{c}}{f_{\text{min}}\cdot T_{sampling}}$,
  $a_{scale}\in[a_{min},a_{max}]$
  $I_{{TFR}}\leftarrow\{\}$
  for ($i_{w},t_{n},T_{n}$) in ($I_{w},t_{w},T_{w}$): do
  Wavelets: $\Gamma_{iw}(a,b)=\int_{-\infty}^{\infty}i_{w}\psi^{*}\left(\frac{t-b}{a}\right)dt.$
  Energy: $E=\sum_{m=1}^{M}\left|\Gamma_{iw}(a,b)\right|^{2}$
  Dominant frequency: $f_{d}=a_{\text{scale}}\left[\arg\max(E)\right]$.
  Entropy: $h=-\sum_{i=m}^{M}P(i_{w}(t))\log P(i_{w}(t))$.
  Kurtosis: $K=\frac{\mathbb{E}[(i_{w}(t)-\mu)^{4}]}{\sigma^{4}}$.
  Skewness: $s_{k}=\frac{\mathbb{E}[(i_{w}(t)-\mu)^{3}]}{\sigma^{3}}$.
  mean: $\mu=\frac{1}{M}\sum_{m=1}^{M}i_{w}(m)$.
  standard deviation: $\sigma=\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left(i_{w}(m)-\mu\right)^{2}}$.
  $X_{n}\leftarrow[\log(E),f_{d},h,K,sk,\mu,\sigma]$
  end for
  return $I_{TFR}=Concat(X_{1},X_{2}\ldots X_{N_{s}},t_{n},T_{n})$
Dataset Explanation and Curation: PRONOSTIA dataset is a widely recognized benchmark dataset in the field of condition monitoring and degradation estimation of rolling-element bearings. Nectiux et al.(nectoux2012pronostia) developed this dataset as part of the PRONOSTIA experimental platform. The dataset comprises 16 complete run-to-failure experiments performed under accelerated wear conditions across three different operating settings: 1800 rpm with 4 kN radial load, 1650 rpm with a 4.2 kN load, and 1500 rpm with a 5 kN load, all at a frequency of 100 Hz. Vibration data were recorded using accelerometers placed along the horizontal and vertical axes, which were sampled at 25.6 kHz. Additionally, temperature readings were collected at a sampling rate of 10 Hz. The data distributions for training and testing are provided in Table 5.
Table 5: Data distributions of PRONOSTIA, XJTU-SY, and HUST datasets.
Dataset Condition Frequency Radial Load Speed Train Test
PRONOSTIA Condition 1 100 Hz 4 kN 1800 rpm $4\sim 7$ $1\sim 3$
Condition 2 100 Hz 4.2 kN 1650 rpm - $1\sim 7$
Condition 3 100 Hz 5 kN 1500 rpm - $1\sim 3$
XJTU-SY Condition 1 35 Hz 4 kN 2100 rpm - $1\sim 5$
Condition 2 37.5 Hz 4.2 kN 2250 rpm - $1\sim 5$
Condition 3 40 Hz 5 kN 2400 rpm - $1\sim 5$
HUST Condition 1 - 0 W - - $1\sim 5$
Condition 2 - 200 W - - $1\sim 5$
Condition 3 - 400 W - - $1\sim 5$
Note: The PRONOSTIA dataset is utilized for training and generalization testing, while the XJTU-SY and HUST datasets are employed to evaluate cross-validation testing. (-) values are either not available or not utilized.
XJTU-SY Dataset is another widely recognized benchmark dataset developed through collaboration between Xi’an Jiaotong University and Changxing Sumyoung Technology (wang2018xjtu). The dataset comprises 15 complete run-to-failure experiments performed under accelerated degradation conditions with three distinct operational settings: 1200 rpm (35 Hz) with a 12 kN radial load, 2250 rpm (37.5 Hz) with an 11 kN radial load, and 2400 rpm (40 Hz) with a 10 kN radial load. Vibrational signals were recorded using an accelerometer mounted on the horizontal and vertical axes and sampled at 25.6 kHz. This dataset is only used for the cross-validation test.
HUST Dataset is a practical dataset developed by Hanoi University of Science and Technology to support research on ball bearing fault diagnosis (hong2023hust). The dataset includes vibration data collected from five bearing types (6204, 6205, 6206, 6207, and 6208) under three different load conditions: 0 W, 200 W, and 400 W. Six fault categories were introduced, consisting of single faults (inner race, outer race, and ball) and compound faults (inner–outer, inner–ball, and outer–ball). Faults were created as early-stage defects in the form of 0.2 mm micro-cracks, simulating real degradation scenarios. The vibration signals were sampled at 51.2 kHz with approximately 10-second recordings for each case. This dataset is only used for the cross-validation test.
Preprocess: Condition monitoring data comprises 1D non-stationary vibrational signals collected from multiple sensors. To extract meaningful information, these signals must be transformed into features that possess meaningful physical interpretability. We utilized the preprocessing proposed in (razzaq2025carle) and labels are generated according to (razzaq2025developingdistanceawareuncertaintyquantification). Initially, the signal is segmented into small, rectangularized vectors using a windowing technique ($w$), enabling better localization of transient characteristics. The continuous wavelet transform (CWT) (aguiar2014continuous) with the Morlet wavelet (lin2000feature) as the mother wavelet is then applied to obtain a time-frequency representation (TFR). The CWT is defined as $\Gamma(a,b)=\int_{-\infty}^{\infty}x_{w}(t)\frac{1}{\sqrt{a}}\psi^{*}\left(\frac{t-b}{a}\right)\,dt$, where $a$ and $b$ denote the scale and translation parameters, respectively, and $\psi$ is the Morlet wavelet function. From the resulting TFR, a compact set of statistical and domain-specific features is extracted to characterize the operational condition of the bearing. The complete feature extraction procedure is described in Algorithm 3.
Neural Network Architecture: The objective of this problem is to design a compact neural network that can effectively model degradation dynamics while remaining feasible for deployment on resource-constrained devices, enabling localized and personalized prognostics for individual machines. To achieve this, we combine a compact convolutional network with NAC. The CNN component extracts spatial degradation features from the training data, while NAC performs temporal filtering to emphasize informative features. This architecture maintains a small model size without sacrificing representational capacity. Full hyperparameter configurations are reported in Table 4.
Evaluation Metric: Score is a metric specifically designed for RUL estimation in the IEEE PHM (nectoux2012pronostia) to score the estimates. The scoring function is asymmetric and penalizes overestimations more heavily than early predictions. This reflects practical considerations, as late maintenance prediction can lead to unexpected failures with more severe consequences than early intervention can.
$\textit{Score}=\sum_{i:\hat{y}_{i}<y_{i}}\left(e^{-\frac{\hat{y}_{i}-y_{i}}{13}}-1\right)+\sum_{i:\hat{y}_{i}\geq y_{i}}\left(e^{\frac{\hat{y}_{i}-y_{i}}{10}}-1\right)$ (48)